OpenStudy (anonymous):

to rent a certain meeting room, a college charges a reservation fee of $49 and an additional fee of $8.90 per hour. the history club wants to spend less than $129.10 on renting the meeting room. what are the possible amounts of time for which they could rent the meeting room? use T for the number of hours the meeting room in rented and solve your inequality for t. i think the answer is T is greater than or equal to 9 , but i always get the signs messed up :( please help!

OpenStudy (jdoe0001):

how did you get \(\bf t \ge 9\) anyway?

OpenStudy (anonymous):

$129.10 - 49 = 80.1 / $8.90 = 9

OpenStudy (jdoe0001):

well.... I'd rather put it this way they want to spend LESS THAN $129.10 there's a base charge, reservation, of a flat fee of $49 so they have to pay that just to try to get it at all per hour, they charge is $8.90, so \(\begin{array}{llll} hour&charge \\\hline\\ 1&8.9(1)\\ 2&8.9(2)\\ 3&8.9(3)\\ ...&....\\ t&8.9t \end{array}\) so they'd end up spending "t hours" and after "t hours" they'd owe \(\bf 49+8.9t\) and they want THAT to be LESS THAN $129.10 so \(\bf 49+8.9t>129.1\implies 8.9t>129.1-49\implies 8.9t>80.1 \\ \quad \\ t>\cfrac{80.1}{8.9}\implies t>9\ hours\)

OpenStudy (anonymous):

the correct answer was t<9